# Northeast Materials Database (NEMAD): Enabling Discovery of High Transition Temperature Magnetic Compounds

### v1

@@ METHODS
[…]
@@
For the structure-informed model, we created two additional features from the structural details. We constructed a crystal system feature for compounds for which the crystal system is available. There exist seven crystal systems, and so this categorical feature was numerically encoded by assigning a different number from 1 to 7 for each of them. From the space groups of the compounds, we then generated the space group feature. There are 230 space groups, and we transform this categorical feature into a numerical representation by assigning different integer numbers between 1 and 230 for each unique space group. Information on the space group provides information about the symmetry of a crystal and its atomic arrangement, which could be useful for the prediction of magnetic transition temperatures or any other magnetic properties. All features along with their corresponding dimensions are provided in the Supplementary Materials.
@@
To classify the materials as nonmagnetic (NM), ferromagnetic (FM), or antiferromagnetic (AFM), we used the random forest classifier. Random Forest is an ensemble learning method that constructs many decision trees during training and generates the class that is a mode of the classes of the individual trees. It is robust against over-fitting and capable of handling high-dimensional feature spaces.
Our database contains only magnetic materials, so it is impossible to make a model only with our data set to classify magnetic and nonmagnetic materials. We took 11,389 stable non-magnetic materials from the Material Project Database[10] and combined them with our database, resulting in 28,168 unique entries for model development. We generated features from chemical composition, as explained in the feature engineering section. We constructed a target feature with available material type information. Since there are three types of materials, we encode this categorical feature numerically by assigning a unique number from 1 to 3 to each material type. Using the Scikit-learn library [54], we randomly divided the dataset into three mutually exclusive partitions: 60% for training, 20% for validation, and the remaining 20% for testing purposes. Most machine learning models include one or more hyperparameters, which cannot be learned from the data and therefore need to be specified beforehand. We specified different sets of hyperparameters in the models, and the validation set later determined the best-performing set. A random forest classifier model was trained, and a five-fold cross-validation technique was used to find the best-performing hyperparameters by initializing GridSearchCV. We then tested our trained model on test set data. The best-performing hyperparameters are as follows: [Number of estimators: 500, Maximum depth: None, Maximum samples split: 2, Minimum samples leaf: 1, Maximum features: sqrt]
@@
In this section, we built several machine learning models such as random forest (RF) regressors, ensemble neural networks (ENN), and extreme gradient boosting (XGBoost) to estimate the Curie and Néel temperatures of corresponding ferromagnetic and antiferromagnetic materials. First, we made a different dataset for both ferromagnetic and anti-ferromagnetic compounds. We have collected compounds from our NEMAD database. A ferromagnetic dataset includes 11923 compounds with different chemical compositions, and an anti-ferromagnetic dataset contains 5389 unique compounds. The distribution of data in the ferromagnetic dataset is not uniform. We have almost 51% of the total compounds with Curie temperature lower than 300K and 22% of the total compounds with Curie temperature greater than 600K. We believe the dataset with this distribution is less appropriate for accurately predicting high Curie temperatures. We created a balanced dataset to address this issue by randomly removing some entries with Curie temperature values lower than 300K. The balanced dataset contains 8249 unique compounds. We then created another data set for the structure-informed model of ferromagnetic compounds. This dataset contains 5,176 entries with additional features like crystal system and space group. Then, we selected the features and targets for the machine learning models. For the Curie temperature prediction model, the target is the value of the Curie temperature, while for the Néel temperature prediction model, the target is the value of the Néel temperature. Both targets contain values in different units. We converted all temperature records to the Kelvin scale. Some of the compounds contain temperatures in a certain range. To get a single value, we took a mean of the two numbers: the initial and final numbers in the range.
Using the Scikit-learn library, we built machine learning models for all datasets mentioned above: the original dataset, balanced dataset, and structure including datasets of ferromagnetic compounds and a dataset of antiferromagnetic compounds. We built a Random Forest model with a 60/20/20 split for training, validation, and testing. The model’s performance was optimized using five-fold cross-validation with the following hyperparameters: [Number of estimators: 1000, Maximum depth: None, samples split: 2, samples leaf: 1, Maximum features: sqrt]. Similarly, we built an XGBoost model with the following hyperparameter: [Number of estimators: 500, Maximum depth: 12, Learning rate: 0.08, Subsample: 0.8, Column sample by tree: 0.6]
We created an ensemble of 30 fully connected neural networks, each with 8 hidden layers. We provided the same input features as fed into the RF and XGBoost models to the input layers of each network. We used the ReLU activation function and the Adam optimizer with a learning rate of 0.001 for all networks. The model was trained on 80% of the dataset, and the remaining 20% was used for testing purposes. The final prediction was the average output from all 30 networks, which outperformed a single neural network.
Evaluation Metrics
For the evaluation of the classification model, we used four metrics such as precision, recall, f1-score of each class of materials and the accuracy of a model[55].
  • Precision quantifies the degree of accuracy of the model’s positive predictions. For each class, it’s the ratio between correct prediction and all instances predicted to belong to that class. High accuracy implies that the model has a low false positive rate. Mathematically,
[…]
  • The model’s recall, a key measure, quantifies its capacity to identify every good case. For each class, recall is the ratio between correct predictions and all actual instances of that class. High recall implies that the model has a low false negative rate. Mathematically,
[…]
  • The F1-score is calculated by taking the harmonic mean of precision and recall. It should balance both metrics in principle. It is useful for unevenly distributed classes. A high F1 score usually indicates that the model performs very well in terms of both precision and recall. Mathematically,
[…]
  • Accuracy is the ratio between correct predictions (TP and TN) among the total number of cases classified. It gives an overall estimate of how often the model is correct across all classes. Mathematically,
    Accuracy=Correct Predictions Total $\text{Accuracy}=\frac{\text{Correct Predictions }}{\text{Total Predictions}}$ (9)
For the evaluation of the regression model, we used three metrics such as coefficient of determination (R2), mean absolute error (MAE), and root mean square Error (RMSE).
  • R2 assesses how effectively a machine learning model explains variance in the target variable compared to a simple mean model. The value ranges from 0 to 1, and higher values generally indicate better fits of the model. Mathematically,
    $R^{2}=1-\frac{\sum_{j=1}^{n}(y_{j}-\hat{y}_{j})^{2}}{\sum_{j=1}^{n}(y_{j}-\bar{y})^{2}}$ (10)
    where $y_{j}$, $\hat{y}_{j}$ and $\bar{y}$ are the actual value, predicted value and mean of the actual value.
  • MAE is the average of the absolute differences between predictions and actual values. It gives an idea of how far the predictions are from the actual values on average. Lower MAE values indicates better model performance.
    $\text{MAE}=\frac{1}{n}\sum_{j=1}^{n}|y_{j}-\hat{y}_{j}|$ (11)
  • RMSE measures the standard deviation of the residuals or simply prediction errors. Lower RMSE value indicate better model performance. RMSE gives a relatively high weight to large errors, which means it’s more sensitive to outliers than MAE.
    $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{j=1}^{n}(y_{j}-\hat{y}_{j})^{2}}$ (12)
Acknowledgements (not compulsory)
@@
This work was supported by the Office of Ba- sic Energy Sciences, Division of Materials Sciences and Engineering, U.S. Department of Energy, under Award No. DE-SC0020221.
@@
Additional information
To include, in this order: Accession codes (where applicable); Competing interests (mandatory statement).
The corresponding author is responsible for submitting a competing interests statement on behalf of all authors of the paper. This statement must be included in the submitted article file.

### v2

@@ Methods
$\text{L2 Stoichiometry Norm }(L2)=\sqrt{\sum_{i=1}^{84}p_{i}^{2}}$ (4)
 $\text{ Entropy }(\bar{S})=-\sum_{i=1}^{84}p_{i}\log{p_{i}}$ (5)
@@
For the structure-informed model, we created two additional features from the structural details. Since the crystal system is a nominal categorical variable consisting of seven distinct types, assigning arbitrary numeric values (such as integers from 1 to 7) could result in the model incorrectly interpreting these values as having a specific order or hierarchy, which is not the case. To address this, we employed the one-hot encoding technique. One-hot encoding is a method that converts categorical variables into a set of binary variables (0s and 1s), ensuring that the model does not assume any implicit ordering between the categories. Each type of crystal system was represented as a separate binary feature. For example, a crystal system type like “Cubic” would be represented as [1, 0, 0, 0, 0, 0, 0], while another type such as “Tetragonal” would be represented as [0, 1, 0, 0, 0, 0, 0]. This transformation allowed us to represent the categorical data in a numerically stable and unbiased way for the machine learning model.
In the case of the space group feature, which consists of more than 100 distinct categories in our case, using one-hot encoding was not feasible due to the high dimensionality it would introduce. The rest of the feature space has fewer than 100 features, so adding more than 100 binary variables for space groups would significantly increase the complexity of the model. To address this, we applied label encoding. Instead of treating the space group as a categorical variable, we assigned a numerical value to each space group based on the average of the target variable for that group. This approach reduced the space group feature to a single dimension while preserving relevant information, making it computationally efficient and manageable within our model.
@@
To classify materials as nonmagnetic (NM), ferromagnetic (FM), or antiferromagnetic (AFM), we trained and evaluated two machine learning classifiers: a RF classifier and an XGBoost classifier. Random Forest is an ensemble learning technique that constructs multiple decision trees and outputs the majority class as the final prediction. It is known for robustness against overfitting and its ability to model complex, high-dimensional feature spaces. Our classification task required distinguishing among three classes, but our database contained only magnetic materials (FM and AFM). To incorporate nonmagnetic examples, we augmented our dataset with 11,389 stable nonmagnetic compounds from the Materials Project database [10], yielding a combined dataset of 35,037 labeled materials.
Features were generated from the chemical composition of each material, as described in the Feature Engineering section. The target variable Type was encoded numerically: 0 = FM, 1 = AFM, 2 = NM. A stratified 60/20/20 split was performed to divide the data into training, validation, and test sets, ensuring balanced representation of all three classes. For both models, classification pipelines were constructed and optimized using grid search with five fold cross-validation, with accuracy as the evaluation metric.The hyperparameters tuned included the number of estimators, maximum tree depth, minimum samples required to split a node, minimum samples per leaf, and the number of features considered at each split. For the XGBoost model, additional parameters such as learning rate, subsample ratio, and column sampling ratio were also optimized.
The Random Forest classifier was implemented using Scikit-learn [60] with the following best-performing hyperparameters: n_estimators=500, max_depth=None, min_samp._split=2, min_samp._leaf=1, max_features=‘sqrt’. We also used class_weight=‘balanced’ to account for class imbalance.The XGBoost classifier was trained using similar procedures, with the following tuned hyperparameters:
n_estimators=800, max_depth=10, learning_rate=0.1, subsample=0.7, colsample_bytree=0.6.
Both models were evaluated on the validation and test sets using accuracy, confusion matrices, and detailed classification reports. The model achieved similar levels of accuracy and classification performance on the training, validation, and test sets, indicating strong generalization and stability across different data splits.
@@
In this section, we trained and evaluated several machine learning models—Random Forest (RF), Extreme Gradient Boosting (XGBoost), and Ensemble Neural Network (ENN)—to estimate the Curie and Néel temperatures of ferromagnetic and antiferromagnetic compounds. Separate datasets were prepared for each magnetic phase using compounds from our NEMAD database. The ferromagnetic dataset initially contained 24,500 entries, and the antiferromagnetic dataset had 9,500 entries. To ensure each chemical composition appeared only once, we grouped entries by composition. When multiple transition temperatures were reported for the same composition, we computed their mean to obtain a unique transition temperature. This preprocessing step resulted in datasets of 15,577 and 7,893 unique compounds, respectively.
The ferromagnetic dataset exhibited a highly skewed temperature distribution, with approximately 47% of entries having Curie temperatures below 300 K and only 22% above 600 K. To address this imbalance, we adopted a stratified undersampling approach. We manually divided the Curie temperature range into fixed bins and identified the bin with the smallest sample count. For each bin, we randomly sampled a fixed number of compounds equal to this minority count, ensuring balanced representation across bins. The resulting balanced dataset comprised 12,461 ferromagnetic and 6,314 antiferromagnetic compounds. We also prepared a structure-informed dataset containing 5,066 ferromagnetic compounds, which included additional structural features like crystal system and space group. For all models, we used input features derived from the chemical formula, as described in the feature engineering section. In the case of the structure-informed dataset, these features were augmented with structural descriptors.
We used three models (RF, ENN, and XGBoost) on both the original and balanced datasets. For the original dataset, we directly trained all models using a stratified 60/20/20 split into training, validation, and test sets, preserving the distribution of Curie temperatures through bin based stratification. For the balanced dataset, we combined this stratified splitting with undersampling and trained ensemble models to improve generalization and capture predictive uncertainty. Specifically, we constructed multiple balanced sets by randomly undersampling the overrepresented low temperature bins while retaining all high temperature compounds in each set. Each ensemble member was trained on a different balanced set, allowing the models to always see the full set of high temperature data and varying portions of the low temperature data. This ensured that all data points from the original dataset were utilized across the ensemble, preventing overfitting to the dominant low-temperature region while still learning from the complete dataset.
The performance of the models was evaluated using R2, mean absolute error (MAE), and root mean square error (RMSE).
Hyperparameters were optimized using grid search combined with five-fold StratifiedKFold cross-validation. Stratification was based on binned Curie temperature values to preserve distribution across folds. For each hyperparameter configuration, we trained an ensemble of five models on undersampled subsets to ensure balanced representation across temperature ranges. The best hyperparameter configuration was chosen based on the average validation performance, calculated across all cross-validation folds and the ensemble models trained in each fold. The best hyperparameters for Random Forest were: n_estimators=1000, max_depth=None, min_samples_split=2, min_samples_leaf=1, and max_features=’sqrt’.
The optimal hyperparameters for XGBoost were: n_estimators=1200, max_depth=12, learning_rate=0.08, subsample=0.8, and colsample_bytree=0.6.
For the final ensemble, we trained 30 models per algorithm on differently sampled subsets of the full training data and averaged their predictions. The standard deviation across these ensemble outputs was used to generate confidence intervals, providing a measure of uncertainty. This method was applied consistently across the original, balanced, and structure-informed datasets.
In the case of ENN, we first split the dataset into 80% training and 20% test sets using stratified sampling based on binned Curie temperatures. To assess model stability and generalization, we performed five-fold StratifiedKFold cross-validation on the training set. In each fold, we trained an ensemble of five neural networks using stratified undersampling to balance temperature ranges, and averaged their predictions to evaluate fold-level performance.
For final testing, we trained an ensemble of 30 fully connected neural networks, each with 8 hidden layers and ReLU activation, using the Adam optimizer with a learning rate of 0.001. Each model was trained on a different stratified-undersampled subset of the full training data for 500 epochs with a batch size of 64. Final predictions were obtained by averaging the outputs of all ensemble members, while the standard deviation across predictions quantified uncertainty.
Data Availability
The full Northeast Materials Database (NEMAD), including all curated magnetic compounds and their associated properties, is publicly accessible at https://www.nemad.org.
Code Availability
All source code used for the machine learning analysis, together with the processed datasets employed to train and evaluate the models, is available at https://github.com/sumanitani/NEMAD-MagneticML and has been archived on Zenodo (https://doi.org/10.5281/zenodo.17042814) [61].
Acknowledgements
@@
This work was supported by the Office of Basic Energy Sciences, Division of Materials Sciences and Engineering, U.S. Department of Energy, under Award No. DE-SC0020221. The authors thank Houssam Sabri for insightful and fruitful discussions that contributed to this work.
@@
Competing Interests Statement
The authors declare no competing interests.